• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name jwexler

    Log entry time 07:52:09 on June11,2010

    Entry number 325677

    keyword=Shift Summary - Owl 6/11/10

    Shift Summary, 06/11/10

    00:00 - Run 4466 in progress. Right arm still down. Called RC, have instructions to not flip HWP before 10AM wien flip.
    00:45 - MCC called to say the beam is off and they don't know when it'll be back. A conversation with MCC shortly after
    01:40 - Notice heater controls suggest heater draws no current.
    01:45 - MCC calls, ready to give us beam. Move to empty, begin short beam recovery.
    02:10 - Request 50uA.
    02:20 - Minor target vacuum (2*10^-5) alarm goes off. Once at 50uA, resume production runs regardless of vacuum.
    02:40 - Target vacuum stablized at ~1.2*10^-5.
    05:15 - Target flammable gas panel high alarm goes off. Low alarm still going. Doesn't seem physically possible. Still concerning, sent tech-on-call a web-based page.
    07:20 - Howard came in. He'd like to get into the hall as soon as possible for dipole work. Cleared it with Jixie, called MCC.

    Runs:
    4466-4473: 596.4 OK multiplets, 453mC